Mean approximation of periodic functions by generalized Karamata means S. A. Settu, Madras, India

Keywords: Fourier series, Modulus of continuity, [F, d_{n}] means, L^{p}-space AMS Subject classification: 42A10, 42A24 Abstract: We find asymptotically precise upper bound over a subclass of Lebesgue integrable functions for the mean deviation of a function from its [F, d_{n}] means.